OpenStudy (anonymous):

5/6c=7/24 Someone please draw it out. I need to show work. :)

OpenStudy (saifoo.khan):

\[\frac56c=\frac{7}{24}\]

OpenStudy (saifoo.khan):

Cross multiply.

OpenStudy (saifoo.khan):

\[5(c) \times 24 = 6 \times 7\]

OpenStudy (saifoo.khan):

\[5c \times 24 = 6 \times 7\]

OpenStudy (saifoo.khan):

120c = 42 \[c=\frac{42}{120}\] \[c=\frac{7}{20}\]

OpenStudy (anonymous):

Still confused

OpenStudy (saifoo.khan):

where?

OpenStudy (anonymous):

c= all that stuff.

OpenStudy (anonymous):

WHy do we cross muliply?

OpenStudy (saifoo.khan):

that's the easy way. otherwise you have to multiply 6 and 24 on both sides. to cancel 6 & 24 on both sides.

OpenStudy (anonymous):

where are you getting 6/24 from?! o.O